# Dose-aware Diffusion Model for 3D Low-count PET: Multi-institutional Validation with Reader Study and Real Low-dose Data

### v2

@@ II-D PET data acquisitions and reconstructions
All the images were reconstructed using vendors’ software provided by United Imaging Healthcare and Siemens Healthineers. All the reconstruction parameters are the same as those used in clinical settings in the respective hospitals.
III Results
III-A Reader study on three commercial scanners
@@
We performed a reader study to evaluate the network performance. 45 patient studies acquired using three different commercial scanners from three hospitals (15 from each hospital) were randomly selected from the testing dataset and included for the reader study. The three types of commercial scanners are the Siemens mCT scanner, the Siemens Vision Quadra scanner, and the United Imaging uExplorer scanner. Three hospitals are the Yale New Haven Hospital in the US, Shanghai Ruijin Hospital in China, and the University Hospital of Bern in Switzerland.
Low-count images were reconstructed by down-sampling the PET listmode events. For the images acquired on the Siemens Vision Quadra and the United Imaging uExplorer scanners, six low-count levels were available, including 1%, 2%, 5%, 10%, 25%, and 50%. For the images acquired using the Siemens mCT scanner, three low-count levels were available, including 5%, 10%, and 20%. Images were reconstructed using vendors’ software from United Imaging Healthcare and Siemens Healthineers. Details on image reconstruction are provided later in this paper (see Methods).
Three nuclear medicine physicians from the Yale-New Haven Hospital participated in the reader study independently. Five 3D images from each patient were provided to the readers simultaneously. The readers were asked to rank the presented images based on the overall image quality. To simulate a clinical reality, images were presented to readers in DICOM format and they were able to choose different display views and adjust the display scale if necessary.
We chose images from the higher 2 low-count levels for the reader study since they would be more practical to achieve in clinical settings. For the patient studies acquired on the Siemens Vision Quadra and the United Imaging uExplorer scanners, five images are the 25% and 50% low-count images, the two corresponding denoised images using the proposed DDPET-3D method, as well as the 100% full-count image. For the patient studies acquired on the Siemens mCT scanner, five images are the 10% and 20% low-count images, the two corresponding denoised images using the proposed DDPET-3D method, as well as the 100% full-count image.
Reader study results are summarized in Fig. 3. For data acquired using the Siemens mCT scanner, all three readers preferred the 10% and 20% denoised results generated by the DDPET-3D method over the 100% full-count images with overall superior average rankings.
@@
All three readers agreed that the proposed DDPET-3D can generate denoised results that are better or at least comparable to the 100% full-count images in terms of overall image quality across the three types of commercial scanners at different hospitals. Sample denoised images acquired at the Shanghai Ruijin Hospital (United Imaging uExplorer scanner) are presented in Fig. 4. DDPET-3D can generate consistent denoised results across a wide range of low-count levels. DDPET-3D was able to recover fine details in the images (e.g., rib bone in 1% results; myocardium, spine, and femur in 2% results). DDPET-3D also effectively suppressed background noise (e.g., liver in patients shown in Fig. 4), and lesions became clearer in the denoised images (blue arrows in Fig. 4). DDPET-3D effectively removed false hot spots (green arrows in the 25% results) in the images and the true lesions/hot spots are easier to identified. True lesions/hot spots and false hot spots were determined based on the 100% full-count images.
III-B Reader study on real low-dose scans
@@
While rebinning the PET listmode events to generate low-count data to simulate a low-dose setting is a common practice in the literature [16, 11, 12, 6, 7, 8, 9, 10, 13, 14, 15, 17], it may affect the random photon rates. To further demonstrate the clinical potential, 20 real low-dose human studies were acquired using a United Imaging uExplorer scanner with an average injected dose of $27.1\pm 5.4\text{MBq}$ (about 10% of full administered dose). Another reader study was performed to assess the image quality. For a more comprehensive comparison, denoised results generated by the Unified Noise-aware Network (UNN) [16] were included in the reader study. UNN was chosen because: (1) similar to DDPET-3D, it also achieves noise-aware denoising; (2) it was among the top 10 winning methods in the 2022 Ultra Low-dose PET Imaging Challenge held at the 2022 IEEE Medical Imaging Conference  [^1] https://ultra-low-dose-pet.grand-challenge.org/leaderboard/.
@@
Reader study was conducted in the similar way as described previously. Three images were presented to the readers simultaneously and they were asked to rank the images based on their overall image quality. Three images are the real low-dose image, denoised images generated by UNN, and denoised images generated by DDPET-3D. As presented in Fig. 5, all three readers agreed that both DDPET-3D and UNN produced superior denoised results compared with the low-dose images. DDPET-3D produced images with overall better rankings than the UNN with statistical significance ($p<0.05$) observed in reader #1 and #2. Even though DDPET-3D also had a overall better ranking in reader #3, $p=0.20$ was observed.
@@
Sample real low-dose denoised images are presented in Fig. 6. No obvious lesion was identified in this patient. Compared with other images, DDPET-3D effectively suppressed excessive noise (e.g., in the liver) and produced images with clearer organ boundaries (e.g., the aorta wall pointed by the blue arrow in Fig. 6).
III-C Generalizability test in a different hospital
To evaluate the cross-center generalizability of the proposed DDPET-3D, we applied the model trained using the uExplorer data acquired at Shanghai Ruijin Hospital in China to patient studies acquired using the same scanner model at the UC Davis Medical Center in the USA.
In addition to different patient characteristics in different countries, two hospitals also implement different reconstruction parameters, reconstruction voxel sizes, and acquisition protocols (see details in Methods). Images acquired at different hospitals also have varied image contrast due to the differences in scan time and acquisition start time. Shanghai Ruijin Hospital implemented a total of 5 min of scan duration. At UC Davis hospital, they implemented a total of 20 min of scan duration, similar to a typical multi-bed position whole-body scan. Additionally, UC Davis Medical Center also doubled the conventional tracer uptake time from 60 min post-injection to 120 min post-injection for improved tumor to background ratio in oncologic imaging. Shanghai Ruijin Hospital implemented the conventional 60 min tracer post-injection time.
Ten patient studies from the UC Davis Medical Center were included in this experiment. Five low-count levels were generated through listmodel rebinning, including 2.5%, 6.25%, 12.5%, 25%, and 50%. Sample denoised results acquired at the UC Davis Medical Center are presented in Fig. 7. The proposed DDPET-3D showed superior generalizability when directly applied to data acquired at a different hospital. Corresponding quantitative measurements are included in Table I.
Another reader study was conducted to evaluate the overall image quality for the 10 UC Davic patients. Similarly, 25% and 50% low-count and denoised images as well as 100% full-count images were included for reader study. As presented in Fig. 8, even though the DDPET-3D denoised images have similar quantitative measurements to low-count images shown in Table I at 25% and 50%, all three readers unanimously gave superior overall rankings to 25% and 50% denoised results than the 100% full-count images. Since the quantitative values shown in Table I were calculated using 100% full-count as reference, they may not correspond to the clinical performance and physicians’ preferences. The reader study results presented in Fig. 8 demonstrated the superior generalizability of the proposed DDPET-3D comparing to other neural networks in another hospital with different patient population and clinical protocols. This result suggests that DDPET-3D might be transferable to another medical center without further network fine-tuning, though further investigations are needed.
III-D Comparison with other baseline denoising methods
@@
TABLE I: Quantitative assessment for different methods on all the testing patients from United Imaging and Siemens scanners. The measurements were obtained by averaging the values on the testing human studies. The proposed method consistently produced promising denoised results regardless of input count levels. The best results among different low-count levels are marked in red. Note that 674 out of 774 testing patients from the Shanghai Ruijin Hospital do not have 50% count reconstructions. $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively, for each metric.
@@
Input 42.515 / 0.706 / 0.738 47.175 / 0.413 / 0.859 51.734 / 0.245 / 0.939 54.459 / 0.180 / 0.967 57.793 / 0.124 / 0.986 59.805 / 0.103 / 0.995
UNN 50.757 / 0.270 0.946 52.736 / 0.216 / 0.966 54.658 / 0.175 / 0.978 55.763 / 0.155 / 0.983 56.940 / 0.137 / 0.988 57.327 / 0.138 / 0.992
DDPET-3D (proposed) 51.673 / 0.244 / 0.957 53.782 / 0.193 / 0.972 56.035 / 0.151 / 0.981 57.495 / 0.130 / 0.986 59.749 / 0.102 / 0.991 62.516 / 0.080 / 0.997
@@
Input 47.376 / 0.523 / 0.856 50.809 / 0.339 / 0.924 54.490 / 0.220 / 0.967 56.769 / 0.174 / 0.982 59.514 / 0.126 / 0.992 61.594 / 0.100 / 0.996
UNN 55.154 / 0.208 / 0.977 56.823 / 0.170 / 0.985 58.470 / 0.141 / 0.989 58.986 / 0.137 / 0.991 60.527 / 0.114 / 0.994 61.877 / 0.098 / 0.995
@@
Input 44.929 / 0.474 / 0.831 48.040 / 0.331 / 0.900 51.272 / 0.227 / 0.947
UNN 48.975 / 0.291 / 0.933 50.514 / 0.243 / 0.954 51.970 / 0.205 / 0.969
DDPET-3D (proposed) 49.950 / 0.262 / 0.953 51.383 / 0.222 / 0.966 53.056 / 0.184 / 0.976
@@
Input 45.410 / 0.310 / 0.871 49.331 / 0.197 / 0.939 52.178 / 0.141 / 0.969 54.719 / 0.105 / 0.985 56.860 / 0.082 / 0.994
UNN 49.538 / 0.193 / 0.957 51.222 / 0.157 / 0.971 52.192 / 0.140 / 0.978 52.950 / 0.128 / 0.983 53.370 / 0.122 / 0.986
@@
TABLE II: Quantitative assessment for different methods. The measurements were obtained by averaging the values on the testing human studies. The proposed method consistently produced promising denoised results regardless of input count levels. $p<0.01$ was observed in all groups when comparing the proposed method with all other methods (except for 50% SSIM values). The best results among different low-count levels are marked in red. Black voxels were removed for calculations. $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively, for each metric.
Shanghai Ruijin Hospital, United Imaging uExplorer Scanner (20 patients $\times$ 6 = 120 studies)

### v3

@@ III-D PET data acquisitions and reconstructions
All the images were reconstructed using vendor software provided by United Imaging Healthcare and Siemens Healthineers. All the reconstruction parameters are the same as those used in clinical settings in the respective hospitals.
IV Results
IV-A Reader study on three commercial scanners
@@
We performed a reader study to evaluate the network performance. 45 patient studies acquired using three different commercial scanners from three hospitals (15 from each hospital) were randomly selected from the testing dataset and included for the reader study. The three types of commercial scanner are the Siemens mCT scanner, the Siemens Vision Quadra scanner, and the United Imaging uExplorer scanner. Three hospitals are Yale New Haven Hospital in the US, Shanghai Ruijin Hospital in China, and the University Hospital of Bern in Switzerland.
Low-count images were reconstructed by down-sampling PET listmode events. For images acquired on Siemens Vision Quadra and United Imaging uExplorer scanners, six low-count levels were available, including 1%, 2%, 5%, 10%, 25%, and 50%. For images acquired using the Siemens mCT scanner, three low-count levels were available, including 5%, 10%, and 20%. The images were reconstructed using vendor software from United Imaging Healthcare and Siemens Healthineers. Details on image reconstruction are provided later in this paper (see Methods).
Three board-certified nuclear medicine physicians from the Yale-New Haven Hospital participated independently in the study of the readers. Five 3D images of each patient were provided to the readers simultaneously. The readers were asked to rank the presented images based on the overall image quality. To simulate clinical reality, the images were presented to the readers in DICOM format, and they were able to choose different display views and adjust the display scale if necessary.
We chose images from the two higher low-count levels for the reader study, since they would be more practical to achieve in clinical settings. For patient studies acquired on the Siemens Vision Quadra and the United Imaging uExplorer scanners, five images are the 25% and 50% low-count images, the two corresponding denoised images using the proposed DDPET-3D method, as well as the 100% full-count image. For the patient studies acquired on the Siemens mCT scanner, five images are the 10% and 20% low-count images, the two corresponding denoised images using the proposed DDPET-3D method, as well as the 100% full-count image.
The reader study results are summarized in Fig. 3. For data acquired using the Siemens mCT scanner, all three readers preferred the 10% and 20% denoised results generated by the DDPET-3D method over the 100% full-count images with overall superior average rankings.
@@
The three readers agreed that the proposed DDPET-3D can generate denoised results that are better or at least comparable to the 100% full-count images in terms of overall image quality across the three types of commercial scanners at different hospitals. Sample denoised images acquired at the Shanghai Ruijin Hospital (United Imaging uExplorer scanner) are presented in Fig. 4. DDPET-3D can generate consistent denoised results across a wide range of low-count levels. DDPET-3D was able to recover fine details in the images (e.g., rib bone in 1% results; myocardium, spine, and femur in 2% results). DDPET-3D also effectively suppressed background noise (e.g., liver in patients shown in Fig. 4), and the lesions became clearer in the denoised images (blue arrows in Fig. 4). DDPET-3D effectively removed false hot spots (green arrows in the 25% results) in the images, and the true lesions/hot spots are easier to identify. True lesions/hot spots and false hot spots were determined based on the 100% full-count images. Please note that we are not claiming perfect reconstruction at extremely low count levels (e.g. 1% or 2%). These examples are included to illustrate the practical limit and robustness of DDPET-3D under challenging conditions.
IV-B Reader study on real low-dose scans
@@
While rebinning the PET listmode events to generate low-count data to simulate a low-dose setting is a common practice in the literature [16, 11, 12, 6, 7, 8, 9, 10, 13, 14, 15, 17], it may affect the random photon rates. To further demonstrate the clinical potential, 20 real low-dose human studies were acquired using a United Imaging uExplorer scanner with an average injected dose of $27.1\pm 5.4\text{MBq}$ (about 10% of full administered dose). Another reader study was performed to assess the image quality. For a more comprehensive comparison, the denoised results generated by the Unified Noise-aware Network (UNN) [16] were included in the reader study. UNN was chosen because: (1) similar to DDPET-3D, it also achieves noise-aware denoising; (2) it was among the top 10 winning methods in the 2022 Ultra Low-dose PET Imaging Challenge held at the 2022 IEEE Medical Imaging Conference  [^1] https://ultra-low-dose-pet.grand-challenge.org/leaderboard/.
@@
The reader study was conducted in the similar way as described previously. Three images were presented to the readers simultaneously and they were asked to rank the images based on their overall image quality. Three images are the real low-dose image, denoised images generated by UNN, and denoised images generated by DDPET-3D. As presented in Fig. 5, all three readers agreed that both DDPET-3D and UNN produced superior denoised results compared to low-dose images. DDPET-3D produced images with overall better rankings than the UNN with statistical significance ($p<0.05$) observed in reader #1 and #2. Even though DDPET-3D also had a overall better ranking in reader #3, $p=0.20$ was observed.
@@
Sample real low-dose denoised images are presented in Fig. 6. In this patient, no obvious lesion was identified. Compared to other images, DDPET-3D effectively suppressed excessive noise (e.g., in the liver) and produced images with clearer organ boundaries (e.g., the aorta wall pointed by the blue arrow in Fig. 6).
IV-C Comparison with other baseline denoising methods
@@
TABLE I: Quantitative comparison of different methods across all test patients from four hospitals. The reported metrics–PSNR, NRMSE, and SSIM–were computed by averaging values over the entire testing dataset. The proposed method consistently produced promising denoised results regardless of input count levels. The best results among different low-count levels are marked in red. Arrows $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively, for each metric. Asterisks (*) denote statistically significant differences ($p<0.05$) between the comparison methods and the proposed DDPET-3D.
@@
Input 42.515* / 0.706* / 0.738* 47.175* / 0.413* / 0.859* 51.734* / 0.245* / 0.939* 54.459* / 0.180* / 0.967* 57.793* / 0.124* / 0.986* 59.805* / 0.103* / 0.995*
UNN 50.757* / 0.270* / 0.946* 52.736* / 0.216* / 0.966* 54.658* / 0.175* / 0.978* 55.763* / 0.155* / 0.983* 56.940* / 0.137* / 0.988* 57.327* / 0.138* / 0.992*
DDIM 41.608* / 0.760* / 0.868* 41.644* / 0.757* / 0.886* 41.652* / 0.757* / 0.898* 41.659* / 0.756* / 0.903* 41.674* / 0.755* / 0.909* 41.106* / 0.768* / 0.907*
DiffusionMBIR 41.827* / 0.742* / 0.903* 41.795* / 0.744* / 0.912* 41.759* / 0.747* / 0.917* 41.745* / 0.749* / 0.919* 41.729* / 0.750* / 0.922* 41.171* / 0.754* / 0.925*
TPDM 41.736* / 0.749* / 0.887* 41.768* / 0.746* / 0.902* 41.736* / 0.749* / 0.910* 41.690* / 0.753* / 0.910* 41.728* / 0.750* / 0.917* 41.192* / 0.753* / 0.921*
Spach Transformer 48.814* / 0.335* / 0.940* 52.381* / 0.223* / 0.968* 55.615 / 0.154 / 0.979* 57.385 / 0.135* / 0.982* 59.385* / 0.116* / 0.989* 61.995 / 0.075 / 0.997
AMIR 45.504* / 0.502* / 0.930* 46.935* / 0.442* / 0.958* 47.998* / 0.405* / 0.973* 48.573* / 0.390* / 0.980* 49.224* / 0.374* / 0.987* 49.722* / 0.320* / 0.994*
DDPET-3D (proposed) 51.673 / 0.244 / 0.957 53.782 / 0.193 / 0.972 56.035 / 0.151 / 0.981 57.495 / 0.130 / 0.986 59.749 / 0.102 / 0.991 62.516 / 0.071 / 0.997
@@
Input 47.376* / 0.523* / 0.856* 50.809* / 0.339* / 0.924* 54.490* / 0.220* / 0.967* 56.769* / 0.174* / 0.982* 59.514* / 0.126* / 0.992 61.594* / 0.100* / 0.996*
UNN 55.154* / 0.208* / 0.977* 56.823* / 0.170* / 0.985* 58.470* / 0.141* / 0.989* 58.986* / 0.137* / 0.991* 60.527* / 0.114* / 0.994* 61.877* / 0.098* / 0.995*
DDIM 43.639* / 0.765* / 0.888* 43.728* / 0.758* / 0.904* 43.759* / 0.755* / 0.911* 43.784* / 0.753* / 0.913* 43.825* / 0.749* / 0.917* 43.852* / 0.747* / 0.920*
DiffusionMBIR 43.807* / 0.751* / 0.910* 43.826* / 0.750* / 0.919* 43.830* / 0.749* / 0.922* 43.840* / 0.748* / 0.922* 43.862* / 0.747* / 0.924* 43.873* / 0.746* / 0.926*
TPDM 43.852* / 0.749* / 0.905* 43.899* / 0.745* / 0.917* 43.926* / 0.743* / 0.923* 43.951* / 0.741* / 0.925* 43.988* / 0.738* / 0.928* 44.010* / 0.736* / 0.930*
Spach Transformer 53.570* / 0.225* / 0.976* 57.255 / 0.163* / 0.985 58.555 / 0.135 / 0.991 60.017 / 0.116* / 0.992 61.232* / 0.093 / 0.993* 64.156 / 0.076* / 0.996
AMIR 49.389* / 0.419* / 0.973* 49.952* / 0.400* / 0.983* 50.399* / 0.387* / 0.989* 50.647* / 0.381* / 0.992 50.915* / 0.374* / 0.993* 51.119* / 0.370* / 0.996
@@
Input 44.929* / 0.474* / 0.831* 48.040* / 0.331* / 0.900* 51.272* / 0.227* / 0.947*
UNN 48.975* / 0.291* / 0.933* 50.514* / 0.243* / 0.954* 51.970* / 0.205* / 0.969*
DDIM 40.060* / 0.721* / 0.821* 40.138* / 0.715* / 0.835* 40.235* / 0.707* / 0.848*
DiffusionMBIR 40.448* / 0.692* / 0.876* 40.462* / 0.690* / 0.882* 40.515* / 0.687* / 0.889*
TPDM 39.784* / 0.698* / 0.863* 39.797* / 0.695* / 0.871* 39.945* / 0.690* / 0.880*
Spach Transformer 49.788* / 0.275* / 0.950* 51.003* / 0.207* / 0.963* 52.281* / 0.188* / 0.975*
AMIR 45.592* / 0.396* / 0.929* 46.557* / 0.362* / 0.954* 47.374* / 0.337* / 0.971*
DDPET-3D (proposed) 49.950 / 0.262 / 0.953 51.383 / 0.199 / 0.966 53.056 / 0.165 / 0.976
@@
Input 45.410* / 0.310* / 0.871* 49.331* / 0.197* / 0.939* 52.178* / 0.141* / 0.969* 54.719 / 0.105 / 0.985 56.860 / 0.082 / 0.994
UNN 49.538* / 0.193* / 0.957* 51.222* / 0.157* / 0.971 52.192* / 0.140* / 0.978 52.950* / 0.128* / 0.983 53.370* / 0.122* / 0.986*
DDIM 37.436* / 0.754* / 0.840* 37.542* / 0.746* / 0.864* 37.578* / 0.743* / 0.874* 37.608* / 0.740* / 0.882* 37.633* / 0.738* / 0.889*
DiffusionMBIR 37.583* / 0.742* / 0.877* 37.636* / 0.738* / 0.891* 37.656* / 0.737* / 0.897* 37.668* / 0.736* / 0.902* 37.674* / 0.735* / 0.905*
TPDM 37.388* / 0.758* / 0.846* 37.551* / 0.745* / 0.871* 37.635* / 0.738* / 0.887* 37.621* / 0.739* / 0.888* 37.494* / 0.750* / 0.876*
Spach Transformer 50.129 / 0.182 / 0.959* 50.953* / 0.150* / 0.968* 52.819* / 0.134* / 0.974* 54.49 / 0.126* / 0.983 55.947 / 0.093* / 0.989
AMIR 43.900* / 0.379* / 0.946* 44.800* / 0.350* / 0.963* 45.475* / 0.333* / 0.971* 46.130* / 0.318* / 0.982* 46.798* / 0.305* / 0.989
@@
Although most existing deep learning methods are not directly applicable to PET image denoising across varying low-count levels, we believe comparisons with other related methods are still beneficial to show the effectiveness of the proposed method. In this subsection, the proposed method is compared with the following methods:
  1. 1.
    Unified Noise-aware Network (UNN)[16]: Similar to the proposed DDPET-3D, UNN was designed for noise-aware PET image denoising. It was among the top 10 winning methods in the 2022 Ultra Low-dose PET Imaging Challenge at the IEEE Medical Imaging Conference, serving as a strong U-Net baseline.
  2. 2.
    DDIM [27]: A straightforward conditional diffusion model was included for comparison. To maintain a sampling speed similar to that of other diffusion models, DDIM sampling was implemented.
  3. 3.
    DiffusionMBIR [38]: A diffusion-based model also aims to address the 3D inconsistency issue in diffusion model by applying a TV penalty along the Z-axis.
  4. 4.
    TPDM [39]: Another diffusion-based method that addresses 3D inconsistency by jointly sampling from two perpendicular diffusion models.
  5. 5.
    Spach Transformer [35]: To capture long-range information, the Spach Transformer implements a transformer-based approach to achieve PET image denoising.
  6. 6.
    AMIR [36]: AMIR is an all-in-one network capable of handling multiple medical imaging tasks (including PET image denoising) using a single unified model.
As presented in Fig. 7, standard DDIM sampling produced images with severe inconsistencies between slices in the sagittal view. Both DiffusionMBIR and TPMD improved the results, but the problem still exists. Although AMIR is not a diffusion-based method, it also resulted in inconsistent reconstruction since it is a 2D-based method. The proposed DDPET-3D method produced noticeably more consistent reconstructions. Moreover, even though other diffusion models produced visually realistic PET images, they failed to accurately reconstruct the structure of different organs. For example, all other diffusion models produced images with distorted myocardium as shown in the transverse slice in Fig. 7. UNN, Spach Transformer, and AMIR also failed to accurately recover the shape of the myocardium. In addition, the proposed DDPET-3D method was also able to recover some fine details in the images, such as the hot spot pointed by the blue arrows in Fig. 7, which is either over-smoothed (in UNN, Spach Transformer, and AMIR) or barely visible (in other diffusion methods). Although DDPET-3D is not perfect, it consistently outperforms the comparison methods in terms of preserving the anatomical structure and recovering clinically relevant details.
Another major issue of the diffusion model is inaccurate image quantification. Note that all the presented images were already normalized by the total injected activities. As shown in Fig. 7, the tracer activities in certain organs are completely wrong in other existing diffusion models. For example, brain activities are noticeably lower in the DDIM, DiffusionMBIR, and TPDM results. Such differences may affect certain diagnostic tasks such as lesion detection [5]. Even though UNN produced over-smoothed reconstructions, it does not alter the overall tracer activities in different organs. Using UNN output as denoised prior, the proposed DDPET-3D maintained overall image quantification. Also, using UNN as a denoised prior allows the proposed method to recover some subtle features that are almost invisible in the low-count input (blue arrows in Fig. 7).
Image quality was quantitatively evaluated using PSNR (peak signal-to-noise ratio), NRMSE (normalized root mean square error), and SSIM (structure similarity index). Quantitative results are presented in Table I.
As presented in Table I, the proposed method outperformed other methods in most of the six low-count levels in the four hospitals. Although performance improved gradually for all methods with increasing count levels, only the proposed DDPET-3D method consistently produced optimal results. Both the proposed method and UNN are aware of noise, but UNN performed worse than input at higher count levels (25% and 50%). Spach Transformer also showed superior quantitative performance at higher count levels, performing comparably to DDPET-3D in certain settings. Also, due to inaccurate image quantification and slice inconsistency, other comparison diffusion models produced images with even worse quantitative results compared to low-count inputs at higher count levels. Additionally, since the AMIR is a 2D method that reconstructs the 3D volume slice-by-slice, it achieved good SSIM performance but showed poor voxel-wise quantification (PSNR and NRMSE) due to slice inconsistency.
IV-D Generalizability test in a different hospital
TABLE II: Quantitative results of the generalizability experiment. The reported metrics–PSNR, NRMSE, and SSIM–were computed by averaging values over the entire testing dataset. Arrows $\uparrow$ and $\downarrow$ indicate that the higher and lower values are better, respectively, for each metric. Asterisks (*) mark statistically significant differences ($p<0.05$) between the model trained on Ruijin data and the model trained on data from the respective hospital. For ease of comparison, some statistics are copied from Table I.